OpenStudy (anonymous):

Stumped. What is the distance between the origin and the point (8, -17)? If necessary, round your answer to two decimal places.

OpenStudy (johnweldon1993):

You can use the distance formula \[d = \sqrt{(x2-x1)^2 + (y2 - y1)^2} \]

OpenStudy (johnweldon1993):

you have a point (8 , -17) and you have the point of the origin (0,0)

OpenStudy (anonymous):

15 units or 25 units

OpenStudy (zzr0ck3r):

 what is C?

OpenStudy (anonymous):

15??

OpenStudy (johnweldon1993):

Not quite lets do it step by step (0 - 8)² = ? (0 - (-17))² = ? so what is √((-8²) + (-17²))

OpenStudy (anonymous):

20.33 units

OpenStudy (johnweldon1993):

-8² = 64 -17² = 289 64 + 289 = 353 √353 = ...?

OpenStudy (anonymous):

18.7

OpenStudy (johnweldon1993):

rounded to 2 decimal places...

OpenStudy (anonymous):

18.80

OpenStudy (johnweldon1993):

√353 = 18.78829423 ^ | 2 decimal places.... so 18.79 right?

OpenStudy (anonymous):

ohh ok yea

OpenStudy (anonymous):

thanks again